OpenStudy (anonymous):

A recipe for 12 biscuits use 200 g of margarine, 400 g of sugar, 500 g of flour and 300 g of grdrice. What is the number of biscuits I could make if I had 1kg of each ingredient

OpenStudy (anonymous):

I worked this out so far total ingredient is 1400 g 1kg of each ingredient would be 4000 g 1400 g =12 biscuits 2800 g= 24 biscuits there is still 1200 grams remaining, not enough for another 12 but enough for 6 more So is the answer 24 or 30?

OpenStudy (anonymous):

Help anyone?

OpenStudy (texaschic101):

1400/12 = 4000/x --- (1400 g makes 12 biscuits = 4000 g to x biscuits) since this is a proportion, we cross multiply (1400)(x) = (12)(4000) 1400x = 48,000 x = 48,000/1400 x = 34.28....rounds to 34 biscuits

OpenStudy (anonymous):

I see,better strategy. Thanks!